Hi Jane bought 1/3 as many almond cookies as chocolate cookies.She gave each of her friends 3 almond & 4 chocolate cookies.She then had 3 almond & 24 chocolate cookies left.
How many friends received cookies from her.
How many chocolate cookies did she buy.
Thanks

Answer by ikleyn(52781)

.

 x = almond cookies

3x = chocolate cookies.


From the condition, we have THESE two equations


    x = 3n + 3     (1)    (n = the number of friends)

   3x = 4n + 24    (2)


Multiply equation (1) by 3 (both sides). You will get

    3x = 9n + 9.


Substitute it into equation (2).  You will get

    9n + 9 = 4n + 24.


It implies

    9n - 4n = 24 - 9

      5n    = 15

       n    = 15/5 = 3.


So, there are 3 friends.


By knowing it, CAN YOU answer the last question on your own ?

Solved.

------------

Hello, does your appreciation stretch so far that you will post your "THANKS" to me for my teaching ?